\begin{thebibliography}{99}

\bibitem{AGHSY} E.\:Akin, E.\:Glasner, W.\:Huang, S.\:Shao, X.\:Ye, Sufficient conditions under which a transitive system is chaotic. Ergodic Theory Dyn. Syst. 30 (2010), 1277--1310.

\bibitem{A} D.V.\:Anosov, Geodesic flows on closed Riemann manifolds with negative curvature. Proc. Steklov Inst. Math. 90 (1967), 235 p.

\bibitem{AH} N.\:Aoki, K.\:Hiraide, Topological theory of dynamical systems. Recent advances. North--Holland Mathematical Library, 52. North--Holland Publishing Co., 1994.

\bibitem{AGW} J.\:Auslander, E.\:Glasner, B.\:Weiss, On recurrence in zero dimensional flows. Forum Math. 19 (2007), 107--114.

\bibitem{BHR} F.\:Blanchard, B.\:Host, S.\:Ruette, Asymptotic pairs in positive-entropy systems. Ergodic Theory Dyn. Syst. 22 (2002), 671--686.

\bibitem{B1} R.\:Bowen, Entropy-expansive maps. Trans. Am. Math. Soc. 164  (1972), 323--331.

\bibitem{B2} R.\:Bowen, Equilibrium states and the ergodic theory of Anosov diffeomorphisms. Lecture Notes in Mathematics, 470. Springer, 1975.

\bibitem{GM} S.\:Glasner, D.\:Maon, Rigidity in topological dynamics. Ergodic Theory Dyn. Syst. 9 (1989), 309--320.

\bibitem{Kat} H.\:Kato, Continuum-wise expansive homeomorphisms. Can. J. Math. 45 (1993), 576--598.

\bibitem{LO1} J.\:Li, P.\:Oprocha, Shadowing property, weak mixing and regular recurrence. J. Dyn. Differ. Equations 25 (2013), 1233--1249.

\bibitem{LO2} J.\:Li, P.\:Oprocha, Properties of invariant measures in dynamical systems with the shadowing property. Ergodic Theory Dyn. Syst. 38 (2018), 2257--2294.

\bibitem{L} E.\:Lindenstrauss, Mean dimension, small entropy factors and an embedding theorem. Publ. Math., Inst. Hautes \'{E}tud. Sci. 89 (1999), 227--262.

\bibitem{M} J.\:van Mill, The infinite-dimensional topology of function spaces. North--Holland Mathematical Library, 64. Elsevier, 2001.

\bibitem{Moo} T.K.S.\:Moothathu, Implications of pseudo-orbit tracing property for continuous maps on compacta. Topology Appl. 158 (2011), 2232--2239.

\bibitem{MO} T.K.S.\:Moothathu, P.\:Oprocha, Shadowing, entropy and minimal subsystems. Monatsh. Math. 172 (2013), 357--378.

\bibitem{P} S.Yu.\:Pilyugin, Shadowing in dynamical systems. Lecture Notes in Mathematics, 1706. Springer, 1999.

\bibitem{W} P.\:Walters, An introduction to ergodic theory. Graduate Texts in Mathematics, 79. Springer, 1982.

\end{thebibliography}

\documentclass[11pt]{amsart}
\usepackage[margin=30mm]{geometry}
\usepackage{amsmath,amssymb}
\usepackage{amsthm}
\usepackage{mathrsfs}

\newtheorem{thm}{Theorem}[section]

\newtheorem{lem}{Lemma}[section]

\newtheorem{defi}{Definition}[section]
\newtheorem{ex}{Example}[section]
\newtheorem{rem}{Remark}[section]

\usepackage{enumitem}

\begin{document}

\title{Some consequences of shadowing and entropy-expansiveness}
\author{Noriaki Kawaguchi}
\subjclass[2020]{37B20, 37B40, 37B65}
\keywords{shadowing, h-expansive, uniformly rigid, periodic points, zero-dimensional}
\address{Department of Mathematical and Computing Science, School of Computing, Institute of Science Tokyo, 2-12-1 Ookayama, Meguro-ku, Tokyo 152-8552, Japan}
\email{gknoriaki@gmail.com}

\begin{abstract}
We show that if a continuous self-map of a compact metric space is h-expansive and satisfies the shadowing property, then every non-empty uniformly rigid subset is zero-dimensional, and hence the set of periodic points is also zero-dimensional if it is not empty.
\end{abstract}

\maketitle

\markboth{NORIAKI KAWAGUCHI}{Some consequences of shadowing and entropy-expansiveness}

\section{Introduction}

{\em Shadowing} is an important notion in the qualitative theory of dynamical systems. Originally introduced in the context of hyperbolic differentiable dynamics \cite{A,B2}, it describes the phenomenon where coarse orbits, or {\em pseudo-orbits}, are closely approximated by true orbits. For an overview of shadowing theory, we refer the reader to the monographs \cite{AH,P}. The notion of {\em rigidity}, first introduced in \cite{GM} as uniform recurrence in topological dynamics, has attracted considerable attention. In particular, \cite{AGHSY} provides results on rigid subsets and their relation to chaotic properties in dynamical systems.

In this paper, we show that if a continuous self-map of a compact metric space is h-expansive and satisfies the shadowing property, then every non-empty uniformly rigid subset must be zero-dimensional, and consequently, the set of periodic points is also zero-dimensional if not empty. The implications of shadowing have been explored in various studies, including \cite{LO1,LO2,Moo,MO}. Notably, it has been shown that for a continuous self-map of a compact metric space with the shadowing property, the set of regularly recurrent points is dense within the chain recurrent set.

We begin with the definition of shadowing property. Throughout, $X$ denotes a compact metric space endowed with a metric $d$. 

\begin{defi}
\normalfont
Let $f\colon X\to X$ be a continuous map and let $\xi=(x_i)_{i\ge0}$ be a sequence of points in $X$. For $\delta>0$, $\xi$ is called a {\em $\delta$-pseudo orbit} of $f$ if $d(f(x_i),x_{i+1})\le\delta$ for all $i\ge0$. For $\epsilon>0$, $\xi$ is said to be {\em $\epsilon$-shadowed} by $x\in X$ if $d(f^i(x),x_i)\leq \epsilon$ for all $i\ge 0$. We say that $f$ has the {\em shadowing property} if for any $\epsilon>0$, there is $\delta>0$ such that every $\delta$-pseudo orbit of $f$ is $\epsilon$-shadowed by some point of $X$.
\end{defi}

We recall the definition of h-expansiveness \cite{B1}. Let $f\colon X\to X$ be a continuous map and let $K$ be a subset of $X$. For $n\ge1$ and $r>0$, a subset $E$ of $K$ is said to be {\em $(n,r)$-separated} if 
\[
\max_{0\le i\le n-1}d(f^i(x),f^i(y))>r
\]
for all $x,y\in E$ with $x\ne y$. For $n\ge1$ and $r>0$, let $s_n(f,K,r)$ denote the largest cardinality of an $(n,r)$-separated subset of $K$. We define $h(f,K,r)$ and $h(f,K)$ by
\[
h(f,K,r)=\limsup_{n\to\infty}\frac{1}{n}\log{s_n(f,K,r)}
\]
and $h(f,K)=\lim_{r\to0}h(f,K,r)$. The topological entropy $h_{\rm top}(f)$ of $f$ is defined by $h_{\rm top}(f)=h(f,X)$ (see, e.g., \cite{W} for more details).

\begin{defi}
\normalfont
For a continuous map $f\colon X\to X$, let
\[
\Phi_{\epsilon}(x)=\{y\in X\colon d(f^i(x),f^i(y))\le\epsilon\:\:\text{for all $i\ge0$}\}
\]
for $x\in X$ and $\epsilon>0$; and let
\[
h_f^\ast(\epsilon)=\sup_{x\in X}h(f,\Phi_{\epsilon}(x))
\]
for $\epsilon>0$. We say that $f$ is {\em h-expansive} if $h_f^\ast(\epsilon)=0$ for some $\epsilon>0$. 
\end{defi}

\begin{rem}
\normalfont
If a continuous map $f\colon X\to X$ is h-expansive, then $h_{\rm top}(f)<\infty$.
\end{rem}

We recall the definition of zero-dimensional subsets. For background on the topological theory of dimension, we refer the reader to \cite{M}. 

\begin{defi}
\normalfont
We say that a subset $S$ of $X$ is {\em zero-dimensional} if it is not empty and if for any $x\in S$ and any open subset $U$ of $S$ with $x\in U$, there is a clopen subset $V$ of $S$ such that $x\in V\subset U$.
\end{defi}

\begin{rem}
\normalfont
A subset $S$ of $X$ is said to be {\em totally disconnected} if it is empty or if every connected component of $S$ is a singleton. We know that if a subset $S$ of $X$ is zero-dimensional, then $S$ is totally disconnected; and the converse holds when $S$ is a non-empty closed subset of $X$.
\end{rem}

Let us recall the definition of uniformly rigid subsets \cite{AGHSY}.

\begin{defi}
\normalfont
Given a continuous map $f\colon X\to X$, we say that a subset $S$ of $X$ is {\em uniformly rigid} if for any $\gamma>0$, there is $l>0$ such that
\[
\sup_{x\in S}d(x,f^l(x))\le\gamma.
\]  
\end{defi}

\begin{rem}
\normalfont
Let $f\colon X\to X$ be a continuous map. If a subset $S$ of $X$ is uniformly rigid, then
\begin{itemize}
\item every subset of $S$ is uniformly rigid,
\item the closure $\overline{S}$ is uniformly rigid.
\end{itemize}
\end{rem}

The main result of this paper is the following theorem.

\begin{thm}
If a continuous map $f\colon X\to X$ is h-expansive and has the shadowing property, then every non-empty uniformly rigid subset $S$ of $X$ is zero-dimensional.
\end{thm}

The following lemma is a special case of the so-called {\em countable closed sum theorem} (see Theorem 3.2.8 of \cite{M}).

\begin{lem}
For any sequence $A_j$, $j\ge1$, of closed subsets of $X$, if $A_j$ is empty or zero-dimensional for all $j\ge1$, then $\bigcup_{j\ge1}A_j$ is either empty or zero-dimensional.
\end{lem}

Given a continuous map $f\colon X\to X$, we denote by $Per(f)$ the set of periodic points for $f$:
\[
Per(f)=\bigcup_{j\ge1}\{x\in X\colon f^j(x)=x\}.
\]
By Theorem 1.1 and Lemma 1.1, we obtain the following theorem.

\begin{thm}
If a continuous map $f\colon X\to X$ is h-expansive and has the shadowing property, then $Per(f)$ is either empty or zero-dimensional.
\end{thm}

\begin{proof}[Proof of Theorem 1.2]
Note that
\[
Fix(f^j)=\{x\in X\colon f^j(x)=x\}
\]
is uniformly rigid for all $j\ge1$. Since $f\colon X\to X$ is h-expansive and has the shadowing property, by Theorem 1.1, $Fix(f^j)$ is empty or zero-dimensional for all $j\ge1$. Since
\[
Per(f)=\bigcup_{j\ge1}Fix(f^j),
\]
and $Fix(f^j)$ is a closed subset of $X$ for all $j\ge1$, by Lemma 1.1, we conclude that $Per(f)$ is either empty or zero-dimensional.
\end{proof}

This paper consists of three sections and an appendix. In the next section, we prove Theorem 1.1. Several examples are given in Section 3. In Appendix A, we prove Theorem A.1. 

\section{Proof of Theorem 1.1}

In this section, we prove Theorem 1.1. For the proof, we need a few lemmas. For a subset $S$ of $X$, we denote by ${\rm diam}\:S$ the diameter of $S$:
\[
{\rm diam}\:S=\sup_{x,y\in S}d(x,y).
\]
Given a continuous map $f\colon X\to X$ and $\delta>0$, a finite sequence $(x_i)_{i=0}^k$ of points in $X$, where $k>0$, is called a {\em $\delta$-chain} of $f$ if $d(f(x_i),x_{i+1})\le\delta$ for all $0\le i\le k-1$.

\begin{lem}
Let $f\colon X\to X$ be a continuous map which is h-expansive and satisfies the shadowing property. Given any $\beta>0$, there is $\gamma>0$ such that for any subset $C$ of $X$, if
\begin{itemize}
\item ${\rm diam}\:C\le\gamma$,
\item $C$ is connected and uniformly rigid,
\end{itemize}
then $\sup_{i\ge0}{\rm diam}\:f^i(C)\le\beta$.
\end{lem}

\begin{proof}
In order to obtain a contradiction, we assume the contrary. Then, there is $\epsilon>0$ such that
\begin{itemize}
\item $h_f^\ast(5\epsilon)=0$,
\item for any $0<\gamma<\epsilon$, there is a subset $C$ of $X$ such that
\begin{itemize}
\item ${\rm diam}\:C\le\gamma$,
\item $C$ is connected and uniformly rigid,
\item ${\rm diam}\:f^k(C)>3\epsilon$ for some $k>0$. 
\end{itemize}
\end{itemize}
We take $\delta>0$ such that every $\delta$-pseudo orbit of $f$ is $\epsilon$-shadowed by some point of $X$. Let $0<\gamma<\epsilon$, $C$, and $k>0$ as above. Since $C$ is uniformly rigid, we have
\[
\sup_{x\in C}d(x,f^l(x))\le\gamma
\]
for some $l>k$. Since $C$ is connected, there are $x,y\in C$ such that
\[
\sup_{0\le i\le l}d(f^i(x),f^i(y))=3\epsilon.
\]
We fix $z\in C$. Let
\[
\gamma_0=(x_i)_{i=0}^l=(z,f(x),f^2(x),\dots,f^{l-1}(x),z)
\]
and
\[
\gamma_1=(y_i)_{i=0}^l=(z,f(y),f^2(y),\dots,f^{l-1}(y),z).
\]
If $\gamma$ is sufficiently small, then $\gamma_0$ and $\gamma_1$ are $\delta$-chains of $f$. Note that
\[
\sup_{0\le i\le l}d(x_i,y_i)=3\epsilon.
\]
For $u=(u_j)_{j\ge1}\in\{0,1\}^\mathbb{N}$, let
\[
\xi_u=(x_i^{u})_{i\ge0}=\gamma_{u_1}\gamma_{u_2}\gamma_{u_3}\cdots,
\]
a $\delta$-pseudo orbit of $f$. Let
\[
Y=\{p\in X\colon\:\text{$\xi_u$ is $\epsilon$-shadowed by $p$ for some $u\in\{0,1\}^\mathbb{N}$}\}
\]
and define a map $\pi\colon Y\to\{0,1\}^\mathbb{N}$ so that $\xi_{\pi(p)}$ is $\epsilon$-shadowed by $p$ for all $p\in Y$. It follows that
\begin{itemize}
\item $Y$ is a closed $f^l$-invariant subset of $X$,
\item $\pi$ is a surjective continuous map with $\pi\circ f^l=\sigma\circ\pi$, where $\sigma\colon\{0,1\}^\mathbb{N}\to\{0,1\}^\mathbb{N}$ is the shift map.
\end{itemize}
Note that
\[
\sup_{i\ge0}d(f^i(p),f^i(q))\le5\epsilon
\]
for all $p,q\in Y$. By taking $p\in Y$, we obtain $Y\subset\Phi_{5\epsilon}(p)$ and so
\[
h_f^\ast(5\epsilon)\ge h(f,\Phi_{5\epsilon}(p))\ge h(f,Y)\ge\frac{1}{l}h(f^l,Y)\ge\frac{1}{l}h(\sigma,\{0,1\}^\mathbb{N})=\frac{1}{l}\log{2}>0,
\]
which is a contradiction. Thus, the lemma has been proved.
\end{proof}

Given a continuous map $f\colon X\to X$, we say that a subset $S$ of $X$ is {\em pairwise recurrent} if for any $x,y\in S$ and $\gamma>0$, there is $l>0$ such that
\[
\max\{d(x,f^l(x)),d(y,f^l(y))\}\le\gamma.
\]

\begin{lem}
Let $f\colon X\to X$ be a continuous map which is h-expansive and satisfies the shadowing property. Given $0<\beta<\epsilon$ and non-empty subsets $E,F$ of $X$, if
\begin{itemize}
\item $h_f^\ast(\epsilon)=0$,
\item $E$ is connected,
\item $E\subset\overline{F}$,
\item $F$ is pairwise recurrent, 
\item $\sup_{i\ge0}{\rm diam}\:f^i(F)\le\beta$,
\end{itemize}
then $E$ is a singleton.
\end{lem}

\begin{proof}
In order to obtain a contradiction, we assume the contrary, i.e., $E$ is not a singleton. Let
\[
0<2\gamma<\min\{{\rm diam}\:E,\epsilon-\beta\}.
\]
We take $\delta>0$ such that every $2\delta$-pseudo orbit of $f$ is $\gamma$-shadowed by some point of $X$. Since $E$ is connected and $E\subset\overline{F}$, there are $x_0,x_1,\dots,x_k\in F$ such that
\begin{itemize}
\item $d(x_0,x_k)>2r$,
\item $\sup_{0\le j\le k-1}d(x_j,x_{j+1})\le\delta$.
\end{itemize}
Since $F$ is pairwise recurrent, for every $0\le j\le k$, we have
\[
\max\{d(x_0,f^{l_j}(x_0)),d(x_j,f^{l_j}(x_j))\}\le\delta
\]
for some $l_j>0$. For $0\le j\le k$, let
\[
\alpha_j=(p^{(j)}_i)_{i=0}^{l_j}=(x_0,f(x_0),\dots,f^{l_j-1}(x_0),x_0),
\]
a $\delta$-chain of $f$. For $0\le j\le k-1$, let
\[
\beta_j=(q^{(j)}_i)_{i=0}^{l_j}=(x_j,f(x_j),\dots,f^{l_j-1}(x_j),x_{j+1}),
\]
a $2\delta$-chain of $f$. Since $\sup_{i\ge0}{\rm diam}\:f^i(F)\le\beta$, we see that
\[
\sup_{0\le i\le l_j}d(p^{(j)}_i,q^{(j)}_i)\le\beta
\]
for all $0\le j\le k-1$. For $1\le j\le k$, let
\[
\tilde{\beta}_j=(r^{(j)}_i)_{i=0}^{l_j}=(x_j,f(x_j),\dots,f^{l_j-1}(x_j),x_{j-1}),
\]
a $2\delta$-chain of $f$. Since $\sup_{i\ge0}{\rm diam}\:f^i(F)\le\beta$, we see that
\[
\sup_{0\le i\le l_j}d(p^{(j)}_i,r^{(j)}_i)\le\beta
\]
for all $1\le j\le k$. Put
\[
L=l_0+l_k+2\sum_{j=1}^{k-1}l_j.
\]
Let
\[
\gamma_0=(y_i)_{i=0}^{L}=\alpha_0\alpha_1\cdots\alpha_{k-1}\alpha_k\alpha_{k-1}\cdots\alpha_1
\]
and
\[
\gamma_1=(z_i)_{i=0}^{L}=\beta_0\beta_1\cdots\beta_{k-1}\tilde{\beta}_k\tilde{\beta}_{k-1}\cdots\tilde{\beta}_1,
\]
$2\delta$-chains of $f$. Note that
\[
2\gamma<\sup_{0\le i\le L}d(y_i,z_i)\le\beta.
\]
For $u=(u_j)_{j\ge1}\in\{0,1\}^\mathbb{N}$, let
\[
\xi_u=(x_i^{u})_{i\ge0}=\gamma_{u_1}\gamma_{u_2}\gamma_{u_3}\cdots,
\]
a $2\delta$-pseudo orbit of $f$. Let
\[
Y=\{z\in X\colon\:\text{$\xi_u$ is $\gamma$-shadowed by $z$ for some $u\in\{0,1\}^\mathbb{N}$}\}.
\]
Defining a map $\pi\colon Y\to\{0,1\}^\mathbb{N}$ as in the proof of Lemma 2.1, we see that 
\begin{itemize}
\item $Y$ is a closed $f^L$-invariant subset of $X$,
\item $\pi$ is a surjective continuous map with $\pi\circ f^L=\sigma\circ\pi$, where $\sigma\colon\{0,1\}^\mathbb{N}\to\{0,1\}^\mathbb{N}$ is the shift map.
\end{itemize}
Note that
\[
\sup_{i\ge0}d(f^i(z),f^i(w))\le\beta+2\gamma\le\epsilon
\]
for all $z,w\in Y$. By taking $z\in Y$, similarly as in the proof Lemma 2.1, we obtain $Y\subset\Phi_{\epsilon}(z)$ and so
\[
h_f^\ast(\epsilon)\ge h(f,\Phi_{\epsilon}(z))\ge h(f,Y)\ge\frac{1}{L}\log{2}>0,
\]
which is a contradiction. This completes the proof of the lemma.
\end{proof}

The proof of the following lemma is left as an exercise for the reader. 

\begin{lem}
Let $D$ be a non-empty closed connected subset of $X$ which is not a singleton. Then, for any $\gamma>0$, there is a non-empty subset $E$ of $D$ such that
\begin{itemize}
\item $E$ is not a singleton,
\item $E$ is connected,
\item ${\rm diam}\:E\le\gamma$.
\end{itemize}
\end{lem}

By Lemmas 2.1, 2.2, and 2.3, we obtain the following lemma.

\begin{lem}
Let $f\colon X\to X$ be a continuous map which is h-expansive and satisfies the shadowing property. Given any non-empty closed subset $D$ of $X$, if $D$ is connected and uniformly rigid, then $D$ is a singleton.
\end{lem}

\begin{proof}
In order to obtain a contradiction, we assume the contrary, i.e., $D$ is not a singleton. Then, by Lemma 2.1 and Lemma 2.3, there are $0<\beta<\epsilon$ and a non-empty subset $E$ of $D$ such that
\begin{itemize}
\item $h_f^\ast(\epsilon)=0$,
\item $E$ is not a singleton,
\item $E$ is connected and uniformly rigid; and so pairwise recurrent,
\item $\sup_{i\ge0}{\rm diam}\:f^i(E)\le\beta$,
\end{itemize}
which contradicts Lemma 2.2. Thus, the lemma has been proved. 
\end{proof}

By Lemma 2.4, we prove Theorem 1.1.

\begin{proof}[Proof of Theorem 1.1]
Let $S$ be a non-empty uniformly rigid subset of $X$. Then, $\overline{S}$ is a non-empty closed uniformly rigid subset of $X$. If $\overline{S}$ is not totally disconnected, then there is a connected component $D$ of $\overline{S}$ which is not a singleton. It follows that $D$ is a non-empty closed subset of $X$ such that
\begin{itemize}
\item $D$ is not a singleton,
\item $D$ is connected and uniformly rigid,
\end{itemize}
which contradicts Lemma 2.4. This implies that $\overline{S}$ is totally disconnected and so zero-dimensional; therefore, we conclude that $S$ is zero-dimensional, completing the proof of the theorem.
\end{proof}

\section{Examples}

In this section, we give several examples.

\begin{ex}
\normalfont
A homeomorphism $f\colon X\to X$ is said to be {\em expansive} if there is $e>0$ such that
\[
\sup_{i\in\mathbb{Z}}d(f^i(x),f^i(y))\le e
\]
implies $x=y$ for all $x,y\in X$. It is known that every expansive homeomorphism $f\colon X\to X$ is h-expansive (see \cite{B1}). If a homeomorphism $f\colon X\to X$ is expansive, then
\[
Fix(f^j)=\{x\in X\colon f^j(x)=x\}
\]
is a finite set for all $j\ge1$; therefore, $Per(f)$ is a countable set and so zero-dimensional. We also know that if a homeomorphism $f\colon X\to X$ is expansive and has the shadowing property, then $Per(f)$ is dense in the chain recurrent set for $f$ (see Chapter 3 of \cite{AH}).
\end{ex}

\begin{ex}
\normalfont
Let $C$ be the Cantor ternary set in the interval $[0,1]$ and let $f=id_C\colon C\to C$, the identity map. By $h_{\rm top}(f)=0$, $f$ is h-expansive. It is easy to see that $f$ has the shadowing property. Note that
\[
Per(f)=Fix(f)=C
\]
is zero-dimensional but an uncountable set.
\end{ex}

\begin{ex}
\normalfont
Let $I=[0,1]$ and let $f=id_I\colon I\to I$, the identity map. By $h_{\rm top}(f)=0$, $f$ is h-expansive. We easily see that $f$ does not have the shadowing property. Note that
\[
Per(f)=Fix(f)=I
\]
is not zero-dimensional. 
\end{ex}

\begin{ex}
\normalfont
Given a sequence $m=(m_j)_{j\ge1}$ of integers such that $2\le m_1<m_2<\cdots$ and $m_j|m_{j+1}$ for each $j\ge1$, let
\begin{itemize}
\item $X(j)=\{0,1,\dots,m_j-1\}$ (with the discrete topology),
\item
\[
X_m=\{(x_j)_{j\ge1}\in\prod_{j\ge1}X(j)\colon x_j\equiv x_{j+1}\pmod{m_j}\:\:\text{for all $j\ge1$}\},
\]
\item $g_m(x)_j=x_j+1\pmod{m_j}$ for all $x=(x_j)_{j\ge1}\in X_m$ and $j\ge1$.
\end{itemize}
We regard $X_m$ as a subspace of the product space $\prod_{j\ge1}X(j)$. The homeomorphism
\[
g_m\colon X_m\to X_m
\]
is called an {\em odometer} with the periodic structure $m$. Since $g_m$ is {\em equicontinuous} and $X_m$ is totally disconnected, $g_m$ has the shadowing property (see, e.g., Theorem 4 of \cite{Moo}). By $h_{\rm top}(g_m)=0$, $g_m$ is h-expansive. Note that $Per(g_m)=\emptyset$. 
\end{ex}

\begin{ex}
\normalfont
We define a map
\[
g\colon[0,1]\to[0,1]
\]
by $g(y)=y^2$ for all $y\in[0,1]$. Let $X=[0,1]^\mathbb{N}$ and let $f\colon X\to X$ be the map defined by
\[
f(x)_j=g(x_j)
\]
for all $x=(x_j)_{j\ge1}\in X$ and $j\ge1$. Since $h_{\rm top}(g)=0$, we have $h_{\rm top}(f)=0$; therefore, $f$ is h-expansive. Since $g$ has the shadowing property, so does $f$. Note that
\[
Per(f)=Fix(f)=\{0,1\}^\mathbb{N}.
\]
\end{ex}

\begin{ex}
\normalfont
Let $X=[0,1]^\mathbb{Z}$ and let $f\colon X\to X$ be the shift map, i.e.,
\[
f(x)_j=x_{j+1}
\]
for all $x=(x_j)_{j\in\mathbb{Z}}\in X$ and $j\in\mathbb{Z}$. We know that $f$ has the shadowing property (see Theorem 2.3.12 of \cite{AH}). Note that
\[
Fix(f)=\{x=(x_j)_{j\in\mathbb{Z}}\in X\colon x_j=x_{j+1}\:\text{for all $j\in\mathbb{Z}$}\}
\]
is not zero-dimensional; therefore, $Per(f)$ is not zero-dimensional. Note also that $h_{\rm top}(f)=\infty$, in particular, $f$ is not h-expansive. Let $S^1=\{z\in\mathbb{C}\colon|z|=1\}$. Let $\alpha\in(0,1)\setminus\mathbb{Q}$ and let $R_\alpha\colon S^1\to S^1$ be the map defined by $R_\alpha(z)=z\cdot e^{2\pi i\alpha}$ for all $z\in S^1$. We know that there is an injective continuous map
\[
h\colon S^1\to X
\]
such that $h\circ R_\alpha=f\circ h$ (see, e.g., \cite{L}). Since $S^1$ is uniformly rigid with respect to $R_\alpha$, $h(S^1)$ is uniformly rigid with respect to $f$. Note that $h(S^1)$ is homeomorphic to $S^1$ and so is not zero-dimensional. 
\end{ex}

\appendix

\section{}

In this Appendix A, we consider closed invariant subsets and present a corollary of Lemma 2.2. For a continuous map $f\colon X\to X$, we say that a subset $S$ of $X$ is {\em $f$-invariant} if $f(S)\subset S$. We say that a subset $S$ of $X$ is {\em pairwise recurrent} if for any $x,y\in S$ and $\gamma>0$, there is $l>0$ such that
\[
\max\{d(x,f^l(x)),d(y,f^l(y))\}\le\gamma.
\]
Note that if a closed $f$-invariant subset $S$ of $X$ is pairwise recurrent, then $f|_S\colon S\to S$ is a homeomorphism; and as a consequence of Proposition 1 of \cite{BHR}, $h_{\rm top}(f|_S)=0$.

A continuous map $f\colon X\to X$ is said to be {\em positively continuum-wise expansive} \cite{Kat} if there is $c>0$ such that for any non-empty closed connected subset $E$ of $X$ which is not a singleton, we have
\[
{\rm diam}\:f^i(E)>c
\]
for some $i\ge0$. By Theorem 5.8 of \cite{Kat}, we know that if a continuous map $f\colon X\to X$ is positively continuum-wise expansive and satisfies $h_{\rm top}(f)=0$, then $X$ is zero-dimensional.

Given a continuous map $f\colon X\to X$ and a closed $f$-invariant subset $S$ of $X$, we say that $f|_S\colon S\to S$ is {\em equicontinuous} if for any $\epsilon>0$, there is $\delta>0$ such that $d(x,y)\le\delta$ implies
\[
\sup_{i\ge0}d(f^i(x),f^i(y))\le\epsilon
\]
for all $x,y\in S$. It is known that if $f|_S$ is surjective and equicontinuous, then $f|_S$ is a homeomorphism and $f|_S^{-1}$ is also equicontinuous. If $f|_S$ is an equicontinuous homeomorphism, then $f|_S$ is {\em distal}, i.e.,
\[
\inf_{i\in\mathbb{Z}}d(f^i(x),f^i(y))>0
\]
for all distinct $x,y\in S$. It is also known that if $f|_S$ is a distal homeomorphism, then $S$ is pairwise recurrent. By results of \cite{AGW}, we know that whenever $S$ is zero-dimensional, the following conditions are equivalent:
 \begin{itemize}
\item $S$ is pairwise recurrent,
\item $f|_S$ is a distal homeomorphism,
\item $f|_S$ is an equicontinuous homeomorphism.
\end{itemize}

By Lemma 2.2 in Section 2, we obtain the following theorem.

\begin{thm}
Let $f\colon X\to X$ be a continuous map which is h-expansive and satisfies the shadowing property. For any non-empty closed $f$-invariant subset $S$ of $X$, if $S$ is pairwise recurrent (in particular, if $f|_S\colon S\to S$ is a distal homeomorphism), then
\begin{itemize}
\item $S$ is zero-dimensional,
\item $f|_S\colon S\to S$ is an equicontinuous homeomorphism. 
\end{itemize}
\end{thm}

\begin{proof}
Since $S$ is pairwise recurrent, we have $h_{\rm top}(f|_S)=0$. We take $0<\beta<\epsilon$ with
\[
h_f^\ast(\epsilon)=0.
\]
From Lemma 2.2, it follows that for any non-empty connected subset $E$ of $S$ which is not a singleton, we have
\[
{\rm diam}\:f^i(E)>\beta
\]
for some $i\ge0$. This implies that $f|_S$ is positively continuum-wise expansive. We conclude that $S$ is zero-dimensional, and thus $f|_S$ is an equicontinuous homeomorphism.   
\end{proof}

\end{document}